the ratio of the measure of two complementary angles is 7:3. Find the measure of each angle.
Solution:
Let x be the common ratio ,
since we have given that ratio between the complementary angle is 7:3
then the two angle are 7x , and 3x
complementary angle means sum of two angles is 90 degree.
hence
7x + 3x = 90
combining the like terms we get
10x = 90
x = 90/10
x= 9
hence two angles are :
7x = 7(9) = 63 degree
3x = 3 ( 9) = 27 degree
Thanks regards

Found 2 solutions by jim_thompson5910, vidya p:
Answer by jim_thompson5910(35256)

7x+3x = 90

10x = 90

x = 90/10

x = 9

7x = 63

3x = 27


The two angles are 63 degrees and 27 degrees.

Answer by vidya p(12)

the ratio of the measure of two complementary angles is 7:3. Find the measure of each angle.
Solution:
Let x be the common ratio ,
since we have given that ratio between the complementary angle is 7:3
then the two angle are 7x , and 3x
complementary angle means sum of two angles is 90 degree.
hence
7x + 3x = 90
combining the like terms we get
10x = 90
x = 90/10
x= 9
hence two angles are :
7x = 7(9) = 63 degree
3x = 3 ( 9) = 27 degree
Thanks regards